OpenStudy (anonymous):

Write a quadratic equation in the variable (x) having the given numbers as solutions. Type the equation in standard form ax^2+bx+c-0, Solution 9, only solution.

OpenStudy (anonymous):

\[(x-9)^2=0\] then algebra to multiply out and put in standard form

sam (.sam.):

(x-9)^(2)=0 x-9=0 x=9

OpenStudy (anonymous):

No, I got it wrong, it says the correct answer is x^2-18x+81